OpenStudy (anonymous):

Help me solve for X step by step! supplementary angles : 4y=2x+180

OpenStudy (anonymous):

Simplifying 4y = 2x + 180 Reorder the terms: 4y = 180 + 2x Solving 4y = 180 + 2x Solving for variable 'y'. Move all terms containing y to the left, all other terms to the right. Divide each side by '4'. y = 45 + 0.5x Simplifying y = 45 + 0.5x

OpenStudy (anonymous):

@OTAKUSUCKSATMath dude

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

i ASKED MY TEACHER AND YOUR ANSWER WAS WAY OFF!